### v2

# Variation of the centered maximal operator

Abstract.
Let $M$ denote the centered Hardy–Littlewood operator on $\mathbb{R}$. We prove that
${\rm Var}(Mf)\leq{\rm Var}(f)-\frac{1}{2}\big{|}|f(\infty)|-|f(-\infty)|\big{|}$
[…]
Key words and phrases:
Maximal operator, bounded variation.
2020 Mathematics Subject Classification:
Primary 42B25; Secondary 26A45.
1. Introduction
Regularity properties of maximal operators have been investigated in numerous papers and various frameworks. Kinnunen [10] initiated the study of boundedness of the Hardy–Littlewood maximal operator on the Sobolev spaces $W^{1,p}(\mathbb{R}^{n})$ for $p\in(1,\infty)$. Subsequently, [10] was complemented by Kinnunen and Lindqvist [11] to the setting of open subsets of $\mathbb{R}^{n}$. See also Tanaka [16] for the case $p=1$ and $n=1$. Starting from the Euclidean setting and the Hardy–Littlewood maximal operators, both centered and uncentered, the study then encompassed their variants such as fractional maximal operators or maximal operators of convolution type; see [6], [7], [13], [14], and references therein. Also discrete analogues of these operators were considered, and related questions in both settings have been recently studied; see [15], [9], [5], [2], [8].
It was an important observation of Aldaz and Pérez Lázaro [1] that maximal operators can actually improve the regularity of involved functions rather than simply to preserve it. They also pointed out the role to be played by variation when measuring the regularity. In particular, it was proved in [1] that if $f\in{\rm BV}(\mathbb{R})$, then for the uncentered Hardy–Littlewood maximal operator $\widetilde{M}$ the function $\widetilde{M}f$ is absolutely continuous on $\mathbb{R}$ and ${\rm Var}(\widetilde{M}f)\leq{\rm Var}(f)$. For the centered operator $M$, on the other hand, Kurka [12] proved
${\rm Var}(Mf)\leq C{\rm Var}(f),\qquad f\in{\rm BV}(\mathbb{R}),$
with some, quite large, $C>0$. It was conjectured that also in this case $C=1$ is enough, see e.g. [12, Section 1]. In the discrete case, i.e., for $f\in{\rm BV}(\mathbb{Z})$ and the discrete counterpart of $M$, the same was conjectured, see [3, Question B]. It is worth recalling the common truth that the uncentered maximal operator has better regularity behavior than its centered counterpart. Furthermore, the proofs of analogous results are usually much subtle in the centered setting.
It seems that both conjectures, although reasonable and expected to be true, do not take the full advantage of how the limits $f(\pm\infty)$ or $F(\pm\infty)$ for $f\in{\rm BV}(\mathbb{R})$ or $F\in{\rm BV}(\mathbb{Z})$, respectively, determine the limits $Mf(\pm\infty)$ or $MF(\pm\infty)$; see Section 3. We thus propose the following strengthening.
Conjecture $\Delta$.
Suppose $f\in{\rm BV}(\mathbb{R})$. Then
(1.1) ${\rm Var}(Mf)\leq{\rm Var}(f)-\frac{1}{2}\big{|}|f(\infty)|-|f(-\infty)|\big{|}.$
Our first main result, Theorem 1.1, supports this conjecture. We mention that there is no need to formulate separately the discrete counterpart of Conjecture $\Delta$, since by our second main result, Theorem 2.2, inequality (1.1) and its discrete analogue are equivalent. Clearly, verification of either version of the conjecture, continuous or discrete, reduces to consideration of nonnegative functions. We shall prove that a certain further reduction to nonnegative simple functions is available; see Proposition 2.3. Finally, we show that the constants $1$ and $\frac{1}{2}$ in (1.1) are optimal; see Section 4.
Let $\mathbb{K}$ be either $\mathbb{R}$ or $\mathbb{Z}$. For $g\colon\mathbb{K}\to\mathbb{C}$ and any nonempty $E\subseteq\mathbb{K}$ define
${\rm Var}_{E}(g)\coloneqq\sup_{\text{monotone }\phi\colon\mathbb{Z}\to E}\sum_{m\in\mathbb{Z}}\big{|}g\big{(}\phi(m+1)\big{)}-g\big{(}\phi(m)\big{)}\big{|},$
which we call the variation of $g$ over $E$. We write $g\in{\rm BV}(\mathbb{K})$ when the total variation ${\rm Var}(g)\coloneqq{\rm Var}_{\mathbb{K}}(g)$ is finite. For each such $g$ the limits $g(\pm\infty)\coloneqq\lim_{x\to\pm\infty}g(x)$ exist and, when $\mathbb{K}=\mathbb{R}$, for each $x\in\mathbb{R}$ the one-sided limits $g(x^{\pm})\coloneqq\lim_{y\to x^{\pm}}g(y)$ exist as well.
Recall the definition of the centered Hardy–Littlewood maximal operator $M$. We set
$Mg(x)\coloneqq\sup_{r>0}\frac{1}{2r}\int_{x-r}^{x+r}|g|\quad\text{and}\quad MG(n)\coloneqq\sup_{m\geq 0}\frac{1}{2m+1}\sum_{y=n-m}^{n+m}|G(y)|$
for any locally integrable $g\colon\mathbb{R}\to\mathbb{C}$ and $x\in\mathbb{R}$, and for any $G\colon\mathbb{Z}\to\mathbb{C}$ and $n\in\mathbb{Z}$. Of course, if $g\in{\rm BV}(\mathbb{R})$, then $g$ is bounded, hence so is $Mg$. The same is true for $MG$ when $G\in{\rm BV}(\mathbb{Z})$.
Our first main result says that (1.1) holds for a certain special class of functions.
Theorem 1.1.
Let $-\infty<A<B<+\infty$, $a,b\in\mathbb{C}$, $K\in\mathbb{N}\coloneqq\{1,2,\dots\}$, and $\{\alpha_{k}\}_{1}^{K}\subset\mathbb{C}$ be given. For a system $-\infty<A=t_{0}<t_{1}<\dots<t_{K}=B<+\infty$ define
$f(x)\coloneqq a{\mathds{1}_{{(-\infty,A)}}}(x)+\sum_{k=1}^{K}\alpha_{k}{\mathds{1}_{{(t_{k-1},t_{k})}}}(x)+b{\mathds{1}_{{(B,\infty)}}}(x),$
for $x\in\mathbb{R}\setminus\{t_{0},\ldots,t_{K}\}$. Arbitrary values can be set for $f(t_{0})$ and $f(t_{K})$, and also for $f(t_{k})$, $11\leq k\leq K-$ 1, unless both $\alpha_{k}$ and $\alpha_{k+1}$ are different from 00 when we set $f(t_{k})=0$. Then
(1.2) ${\rm Var}(Mf)\leq{\rm Var}(f)-\frac{1}{2}\big{|}|a|-|b|\big{|}.$
Similarly, replacing the system of real numbers $\{t_{k}\}_{0}^{K}$ by a system of integers $-\infty<A=n_{0}<n_{1}<\dots<n_{K}=B<+\infty$, and assuming $\alpha_{k}\alpha_{k+1}=0$, $11\leq k\leq K-$ 1, define
$F(n)\coloneqq a{\mathds{1}_{{(-\infty,A)}}}(n)+\sum_{k=1}^{K}\alpha_{k}{\mathds{1}_{{[n_{k-1},n_{k})}}}(n)+b{\mathds{1}_{{[B,\infty)}}}(n)$
for all $n\in\mathbb{Z}$. Then
(1.3) ${\rm Var}(MF)\leq{\rm Var}(F)-\frac{1}{2}\big{|}|a|-|b|\big{|}.$
Theorem 1.1 implies (1.2) or (1.3) for all indicator functions of bounded variation, $f\colon\mathbb{R}\to\{0,1\}$ or $F\colon\mathbb{Z}\to\{0,1\}$, respectively, thus strengthening [4, Theorems 1.1 and 1.3]. Notice a separation condition encoded in the considered systems $\{t_{k}\}_{0}^{K}$ and $\{n_{k}\}_{0}^{K}$, when we require either $f(t_{k})=0$, in some circumstances depending on the behavior of $f$ around $t_{k}$, or $\alpha_{k}\alpha_{k+1}=0$, respectively.
Throughout the paper, $I$ with possible subscripts affixed will denote an open bounded interval on $\mathbb{R}$. Then $c(I),l(I),r(I)$ will be the center and the left/right endpoint of $I$, respectively. As already mentioned, ${\rm Var}$ with no subscripts, $\mathbb{R}$ or $\mathbb{Z}$, will denote the total variation; it will be clear from the context to which situation this symbol refers to. Appearance of a subscript will indicate that the variation is related to a subset represented by this subscript.
2. Transference
The main goal of this section is to prove a transference result, Theorem 2.2, which seems to be of independent interest. We begin with a simple result that exhibits a substantial improvement of regularity of $M$. Namely, given $g\in{\rm BV}(\mathbb{R})$, not only does $Mg$ have one-sided limits at every $x\in\mathbb{R}$ but it is one-sided continuous as well. Of course, the first claim follows, since $Mg\in{\rm BV}(\mathbb{R})$, cf. [12], but we decided to provide a short proof of this fact to make the whole argument self-contained.
Lemma 2.1.
For a generic function $g\in{\rm BV}(\mathbb{R})$ and $x\in\mathbb{R}$ the limits $Mg(x^{\pm})$ exist and we have
(2.1) $Mg(x)=\min\{Mg(x^{-}),Mg(x^{+})\}.$
Proof.
We can assume $g\geq 0$. Regarding the existence of the limits, fix $\varepsilon>0$ and take $\delta>0$ such that $|g(y)-g(x^{-})|<\varepsilon$ for $y\in(x-2\delta,x)$ and $|g(y)-g(x^{+})|<\varepsilon$ for $y\in(x,x+2\delta)$. Let $\tilde{g}(y)$ be $g(x^{-})$, $g(x^{+})$, or $g(y)$, if $y$ belongs to $(x-2\delta,x)$, $[x,x+2\delta)$, or $\mathbb{R}\setminus(x-2\delta,x+2\delta)$, respectively. Then $|Mg(y)-M\tilde{g}(y)|\leq\varepsilon$ for $y\in\mathbb{R}$. Note that $M\tilde{g}=\max\{M_{<\delta}\tilde{g},M_{\geq\delta}\tilde{g}\}$, where $M_{<\delta}$ and $M_{\geq\delta}$ allow only the centered intervals $I$ with $|I|<2\delta$ and $|I|\geq 2\delta$, respectively. If $y\in(x,x+\delta)$, then
$M_{<\delta}\tilde{g}(y)=\max\Big{\{}g(x^{+}),\frac{(\delta-y+x)g(x^{-})+(\delta+y-x)g(x^{+})}{2\delta}\Big{\}},$
whereas if $x<y<z<x+\delta$, then
$|M_{\geq\delta}\tilde{g}(z)-M_{\geq\delta}\tilde{g}(y)|\leq\frac{(z-y){\rm Var}(g)}{2\delta},$
since for $r\geq\delta$ we have
[…]
Thus, $\limsup_{t\to x^{+}}Mg(t)-\liminf_{t\to x^{+}}Mg(t)\leq 2\varepsilon$. Analogously we show that $Mg(x^{-})$ exists.
Regarding (2.1), we have $Mg(x)\leq\min\{Mg(x^{-}),Mg(x^{+})\}$ by lower semicontinuity of $Mg$. By using $Mg(x)\geq\frac{g(x^{-})+g(x^{+})}{2}$ and $Mg(x^{-})\geq\max\{g(x^{-}),\frac{g(x^{-})+g(x^{+})}{2}\},$ and standard symmetry arguments, we conclude that to show the opposite inequality it suffices to prove the implication
$Mg(x^{-})>\max\Big{\{}g(x^{-}),\frac{g(x^{-})+g(x^{+})}{2}\Big{\}}\implies Mg(x)\geq Mg(x^{-}).$
But if the left part above holds, then there is $\delta>0$ such that $Mg(y)=M_{\geq\delta}g(y)$ for $y\in(x-\delta,x)$. Since like before $|M_{\geq\delta}g(x)-M_{\geq\delta}g(y)|\leq\frac{(x-y){\rm Var}(g)}{2\delta}$ for all such $y$, the proof is concluded. ∎
Now, for $\mathbb{K}=\mathbb{R}$ or $\mathbb{K}=\mathbb{Z}$, and $a,b\in\mathbb{C}$, define
${\rm BV}_{a}^{b}(\mathbb{K})\coloneqq\{g\in{\rm BV}(\mathbb{K}):g(-\infty)=a,\,g(+\infty)=b\}.$
By $c(\mathbb{K})$ we denote the smallest value $c\in(0,\infty]$ such that
${\rm Var}(Mg)\leq c{\rm Var}(g),\qquad g\in{\rm BV}(\mathbb{K}),$
and by $C_{a}^{b}(\mathbb{K})$ we denote the largest value $C\in[0,\infty)$ such that
${\rm Var}(Mg)\leq c(\mathbb{K}){\rm Var}(g)-C,\qquad g\in{\rm BV}_{a}^{b}(\mathbb{K}),$
with the convention $C_{a}^{b}(\mathbb{K})=0$ if $c(\mathbb{K})=\infty$.
The transference result that follows complements [4, Proposition 1.4]. We include the potential case $c(\mathbb{K})=\infty$ to emphasize that the proof works independently of whether $c(\mathbb{K})$ is finite or not. Nonetheless, in [12] it was shown that $c(\mathbb{R})<\infty$, thus Theorem 2.2, even in the weaker form with the additional assumption $c(\mathbb{K})<\infty$, immediately gives $c(\mathbb{Z})<\infty$.
Theorem 2.2.
We have $c(\mathbb{Z})=c(\mathbb{R})$ and $C_{a}^{b}(\mathbb{Z})=C_{a}^{b}(\mathbb{R})$ for all $a,b\in\mathbb{C}$.
Proof.
We first show $c(\mathbb{Z})\leq c(\mathbb{R})$ and $C_{a}^{b}(\mathbb{Z})\geq C_{a}^{b}(\mathbb{R})$ by using a natural extension procedure, cf. [4, Lemma 4.1]. If $c(\mathbb{R})=\infty$, then $C_{a}^{b}(\mathbb{R})=0$ and there is nothing to show. Assume therefore that $c(\mathbb{R})<\infty$. Given $G\in{\rm BV}_{a}^{b}(\mathbb{Z})$, consider
$g\coloneqq\sum_{n\in\mathbb{Z}}G(n){\mathds{1}_{{[n-1/2,n+1/2)}}}.$
Then ${\rm Var}(g)={\rm Var}(G)$ so that $g\in{\rm BV}_{a}^{b}(\mathbb{R})$. Moreover, $Mg(n)=MG(n)$ for $n\in\mathbb{Z}$, since in this case the averages of $|g|$ over the centered intervals $I$ of length $|I|\in 2\mathbb{N}-1$ dominate the remaining ones. Thus, ${\rm Var}(MG)={\rm Var}_{\mathbb{Z}}(Mg)\leq{\rm Var}(Mg)$ and the claimed inequalities follow.
In the opposite direction, we claim that $c(\mathbb{Z})\geq c(\mathbb{R})$ and $C_{a}^{b}(\mathbb{Z})\leq C_{a}^{b}(\mathbb{R})$, and use a convenient variant of a sampling procedure. Obviously, we can assume that $c(\mathbb{Z})<\infty$. Given $g\in{\rm BV}_{a}^{b}(\mathbb{R})$ and $\varepsilon>0$, in view of (2.1), there is $N_{*}=N_{*}(\varepsilon)\in\mathbb{N}$ such that for $E_{N_{*}}\coloneqq 2^{-N_{*}}\mathbb{Z}\cap[-2^{N_{*}},2^{N_{*}}]$ we have
(2.2) ${\rm Var}_{E_{N_{*}}}(Mg)\geq{\rm Var}(Mg)-\varepsilon\quad\text{or}\quad{\rm Var}_{E_{N_{*}}}(Mg)\geq\varepsilon^{-1},$
depending on whether ${\rm Var}(Mg)$ is finite or not, respectively. For $N\in\mathbb{N}$, let $M_{N}$ be the variant of $M$ allowing only the centered intervals $I$ with $|I|\in(2\mathbb{N}-1)/2^{N}$. We have $\lim_{N\to\infty}M_{N}g=Mg$ pointwise, since for $x\in\mathbb{R}$ and $\delta>0$ there are $0<\rho_{1}<\rho_{2}<\infty$ such that $\frac{1}{2\rho}\int_{x-\rho}^{x+\rho}|g|\geq Mg-\delta$ for all $\rho\in[\rho_{1},\rho_{2}]$. Thus, there exists $N_{0}\in\mathbb{N}$ such that ${\rm Var}_{E_{N_{*}}}(M_{N}g)\geq{\rm Var}_{E_{N_{*}}}(Mg)-\varepsilon$ for all $N\geq N_{0}$. Consider $\tilde{G}_{N}(n)\coloneqq 2^{N}\int_{(n-1/2)/2^{N}}^{(n+1/2)/2^{N}}|g|$ and $G_{N}(n)\coloneqq 2^{N}\int_{(n-1/2)/2^{N}}^{(n+1/2)/2^{N}}g$. Then
(2.3) $\max\{{\rm Var}(\tilde{G}_{N}),{\rm Var}(G_{N})\}\leq\int_{-1/2}^{1/2}{\rm Var}_{(\mathbb{Z}+t)/2^{N}}(g)\,{\rm d}t\leq{\rm Var}(g),$
so that $\tilde{G}_{N}\in{\rm BV}_{|a|}^{|b|}(\mathbb{Z})$ and $G_{N}\in{\rm BV}_{a}^{b}(\mathbb{Z})$. Notice that $MG_{N}(n)\leq M\tilde{G}_{N}(n)=M_{N}g(n/2^{N})$ for all $n\in\mathbb{N}$. Now, we show that taking $N\geq\max\{N_{*},N_{0}\}$ large enough we obtain $MG_{N}(n)\geq M\tilde{G}_{N}(n)-\varepsilon/(10N_{*}^{2})$ for all $n\in E_{N_{*}}^{N}\coloneqq\{2^{N}Q:Q\in E_{N_{*}}\}\subset\mathbb{N}$. Given $Q\in E_{N_{*}}$, we have
$|I_{Q}|^{-1}\int_{I_{Q}}|g|\geq M\tilde{G}_{N}(2^{N}Q)-\varepsilon/(20N_{*}^{2})$
for some $I_{Q}$ centered at $Q$. For each $N$ large enough, let $I_{Q}^{N}$ be the longest interval centered at $Q$, contained in $I_{Q}$, with $|I_{Q}^{N}|\in(2\mathbb{N}-1)/2^{N}$. Denoting $g_{N}\coloneqq\sum_{n\in\mathbb{Z}}G_{N}(n){\mathds{1}_{{[(n-1/2)/2^{N},(n+1/2)/2^{N})}}}$, we see that $\lim_{N\to\infty}|I_{Q}^{N}|^{-1}\int_{I_{Q}^{N}}|g_{N}|=|I_{Q}|^{-1}\int_{I_{Q}}|g|$ by the dominated convergence theorem on $I_{Q}$, because $g$ is continuous almost everywhere and bounded. Since $MG_{N}(2^{N}Q)\geq|I_{Q}^{N}|^{-1}\int_{I_{Q}^{N}}|g_{N}|$ and $E_{N_{*}}$ is finite, the desired bounds follow for some large $N\geq\max\{N_{*},N_{0}\}$. Consequently,
(2.4) ${\rm Var}_{E_{N_{*}}}(Mg)-\varepsilon\leq{\rm Var}_{E_{N_{*}}}(M_{N}g)={\rm Var}_{E_{N_{*}}^{N}}(M\tilde{G}_{N})\leq{\rm Var}_{E_{N_{*}}^{N}}(MG_{N})+\varepsilon\leq{\rm Var}(MG_{N})+\varepsilon.$
Assuming ${\rm Var}(Mg)=\infty$ gives
$\varepsilon^{-1}\leq{\rm Var}_{E_{N_{*}}}(Mg)\leq{\rm Var}(MG_{N})+2\varepsilon\leq c(\mathbb{Z}){\rm Var}(g)-C_{a}^{b}(\mathbb{Z})+2\varepsilon,$
which contradicts our assumption $c(\mathbb{Z})<\infty$. Hence, we have ${\rm Var}(Mg)<\infty$ (this includes the case $c(\mathbb{R})<\infty$), and applying successively the first part of (2.2), (2.4), (2.3) for $G_{N}$, and the definitions of $c(\mathbb{Z})$ and $C_{a}^{b}(\mathbb{Z})$, and then letting $\varepsilon\to 0$, we verify the claim. ∎
Later on, we shall refer to the two methods from the proof of Theorem 2.2 as to the extension procedure and the sampling procedure, respectively.
Below we show that proving Conjecture $\Delta$ reduces to consideration of nonnegative simple functions. We call $G\in{\rm BV}(\mathbb{Z})$ simple, abusing slightly the common terminology, if there exist $a,b\in\mathbb{C}$ and $N\in\mathbb{N}$ such that $G(-n)=a$ and $G(n)=b$ when $n>N$. Similarly, $g\in{\rm BV}(\mathbb{R})$ is simple if $G(-x)=a$ and $G(x)=b$ when $x>N$, and $g$ is constant on the intervals $[n-\frac{1}{2},n+\frac{1}{2})$, $n\in\mathbb{Z}$.
Proposition 2.3.
Conjecture $\Delta$ is true if (1.1) is satisfied by nonnegative simple functions.
Proof.
It is convenient to prove a version of the above for the discrete case, $\mathbb{K}=\mathbb{Z}$, and then apply Theorem 2.2 and the extension procedure to obtain the result for $\mathbb{K}=\mathbb{R}$.
Given nonnegative $G\in{\rm BV}(\mathbb{Z})$ and $\varepsilon>0$, we have ${\rm Var}_{[-N_{*},N_{*}]\cap\mathbb{Z}}(MG)\geq{\rm Var}(MG)-\varepsilon$ for some $N_{*}=N_{*}(\varepsilon)\in\mathbb{N}$. Note that $G\in{\rm BV}_{a}^{b}(\mathbb{Z})$ for some $a,b\in[0,\infty)$. We choose $N=N(\varepsilon,N_{*})\in\mathbb{N}$ such that $|G(-n)-a|<\varepsilon/(5N_{*})$ and $|G(n)-b|<\varepsilon/(5N_{*})$ if $n>N$. Let $G_{N}(n)$ be equal to $a$, $G(n)$, or $b$, for $n<N$, $n\in[-N,N]$, or $n>N$, respectively. Then
${\rm Var}_{[-N_{*},N_{*}]\cap\mathbb{Z}}(MG_{N})\geq{\rm Var}_{[-N_{*},N_{*}]\cap\mathbb{Z}}(MG)-\varepsilon\geq{\rm Var}(MG)-2\varepsilon,$
because $|MG_{N}(n)-MG(n)|\leq\varepsilon/(5N_{*})$. Since ${\rm Var}(G_{N})\leq{\rm Var}(G)$, we are done. ∎
3. Proof of Theorem 1.1
For a generic function $g\in{\rm BV}(\mathbb{K})$ one has
$Mg\geq\frac{|g(-\infty)|+|g(\infty)|}{2}\quad\text{and}\quad Mg(\pm\infty)=\max\Big{\{}|g(\pm\infty)|,\frac{|g(-\infty)|+|g(\infty)|}{2}\Big{\}}.$
We say that $Mg$ has a local maximum at $x\in\mathbb{K}$ if there exist $y^{\prime},y^{\prime\prime}\in\mathbb{K}$ such that $y^{\prime}<x<y^{\prime\prime}$ and
$\max\{Mg(y^{\prime}),Mg(y^{\prime\prime})\}<\sup\{Mg(y):y\in\mathbb{K}\cap[y^{\prime},y^{\prime\prime}]\}=Mg(x);$
a local minimum of $Mg$ at $x$ is defined analogously.
We are ready to prove Theorem 1.1. By the extension procedure it suffices to deal with $\mathbb{K}=\mathbb{R}$. Indeed, given $F\in{\rm BV}_{a}^{b}(\mathbb{Z})$ as in Theorem 1.1 with the corresponding system $\{n_{k}\}_{0}^{K}$, define $f\in{\rm BV}_{a}^{b}(\mathbb{R})$ with the corresponding system $\{t_{k}\}_{0}^{K}$, $t_{k}=n_{k}-\frac{1}{2}$, by $f\coloneqq\sum_{n\in\mathbb{Z}}F(n){\mathds{1}_{{[n-1/2,n+1/2)}}}$. By using the separation condition $\alpha_{k}\alpha_{k+1}=0$ for $F$, we verify that $f$ has the form as in Theorem 1.1. Moreover, ${\rm Var}(f)={\rm Var}(F)$ and ${\rm Var}(Mf)\geq{\rm Var}(MF)$. Then (1.3) for $F$ follows from (1.2) for $f$.
Until the end of this section we fix $f$ as in Theorem 1.1 and denote $I_{k}=(t_{k-1},t_{k})$, $k=1,\ldots,K$, the corresponding intervals. Without any loss of generality we can consider nonnegative $f$ only, hence from now on we assume $a,b\geq 0$ and $\alpha_{k}\geq 0$ for $k=1,\ldots,K$. Clearly, $Mf$ may not have local maxima at all. However, if $x$ is a local maximum of $Mf$, then two cases may occur:
  1. (1)
    $Mf$ is constant on some interval $J\ni x$, i.e., each point of $J$ is also a local maximum of $Mf$;
  2. (2)
    there is no $J$ with this property, i.e., $x$ is isolated.
In the first case, considering $J$ as the longest interval with the declared property, we shall choose exactly one element $x_{*}\in J$ as a representative and this local maximum will be called substantial. In addition every isolated maximum from the second case, if any, will also be called substantial. It follows from Lemma 3.1 that the number of substantial maxima is finite and does not exceed $K^{2}$.
Lemma 3.1.
Let $x$ be a local maximum of $Mf$. Then there exists $x_{*}$ (either equal to $x$ or such that $Mf$ is constant on the closed interval connecting $x$ and $x_{*}$) with the property that $Mf(x_{*})$ is attained for some $I_{x_{*}}$ centered at $x_{*}$ with $l(I_{x_{*}})=l(I_{k^{\prime}}),r(I_{x_{*}})=r(I_{k^{\prime\prime}})$ for some $1\leq k^{\prime}\leq k^{\prime\prime}\leq K$.
Proof.
Take any sequence $\{\rho_{n}\}\subset(0,\infty)$ such that $Mf(x)=\lim_{n\to\infty}\frac{1}{2\rho_{n}}\int_{x-\rho_{n}}^{x+\rho_{n}}f$. Clearly, if $\{\rho_{n}\}$ has a subsequence diverging to $\infty$, then $Mf(y)\geq Mf(x)$ for all $y\in\mathbb{R}$. Thus, $\{\rho_{n}\}$ is bounded.
Assume $\{\rho_{n}\}$ contains a subsequence converging to 00. Then $Mf(x)=\frac{f(x^{-})+f(x^{+})}{2}$. This implies $f(x^{-})=f(x^{+})=Mf(x)$ by $Mf(x)\geq Mf(x^{\pm})\geq f(x^{\pm})$, where the first inequality holds because $x$ is a local maximum. Note that $x\leq A$ cannot happen because then $Mf(x)=f(x^{-})=a$ so that $Mf(y)\geq Mf(x)$ for all $y\leq x$. Similarly, $x\geq B$ is impossible. Thus, $x\in I_{k}$ and $Mf(x)=\alpha_{k}$ for some $k$. We can replace $x$ by $c(I_{k})$, since $Mf(y)\geq Mf(x)$ for all $y\in I_{k}$ so that $Mf$ is constant on $I_{k}$ because $x$ is a local maximum. Taking $x_{*}=c(I_{k})$ and $I_{x_{*}}=I_{k}$, we obtain the claimed property.
If $\{\rho_{n}\}$ is separated from 00, then it contains a subsequence converging to a positive number. Consequently, $Mf(x)=\frac{1}{|I_{x}|}\int_{I_{x}}f$ for some $I_{x}$ centered at $x$. Note that $f\big{(}l(I_{x})^{+}\big{)}\geq Mf(x)$, since otherwise we could show $Mf(y)>Mf(x)$ for some $\delta>0$ and all $y\in(x,x+\delta)$ by taking $I_{y}$ with $r(I_{y})=r(I_{x})$. Moreover, $l(I_{x})\geq A$ because otherwise we would have $Mf(y)\geq Mf(x)$ for all $y\leq x$. Thus, $l(I_{x})\in\big{[}l(I_{k^{\prime}}),r(I_{k^{\prime}})\big{)}$ for some $k^{\prime}$. Similarly, $f\big{(}r(I_{x})^{-}\big{)}\geq Mf(x)$ and $r(I_{x})\in\big{(}l(I_{k^{\prime\prime}}),r(I_{k^{\prime\prime}})\big{]}$ for some $k^{\prime\prime}\geq k^{\prime}$. Let $x_{*}=\frac{l(I_{k^{\prime}})+r(I_{k^{\prime\prime}})}{2}$. If $x=x_{*}$, then $Mf(x)$ is attained for $\big{(}l(I_{k^{\prime}}),r(I_{k^{\prime\prime}})\big{)}$, as desired. If not, then we can show $Mf(y)\geq Mf(x)$ for all $y$ in the closed interval connecting $x$ and $x_{*}$ by taking all $I$ with $l(I_{k^{\prime}})\leq l(I)\leq l(I_{x})$ and $r(I_{x})\leq r(I)\leq r(I_{k^{\prime\prime}})$. Since $x$ is a local maximum, $Mf$ is constant on this closed interval and $Mf(x_{*})$ is attained for $\big{(}l(I_{k^{\prime}}),r(I_{k^{\prime\prime}})\big{)}$, as desired. ∎
We conclude with noting that each substantial local maximum $x$ of $Mf$, if exists, is of the form $x=\frac{l(I_{k^{\prime}})+r(I_{k^{\prime\prime}})}{2}$ for some $1\leq k^{\prime}\leq k^{\prime\prime}\leq K$, and $Mf(x)=\frac{1}{|I_{x}|}\int_{I_{x}}f$ for $I_{x}=\big{(}l(I_{k^{\prime}},r(I_{k^{\prime\prime}})\big{)}$. Moreover, if $Mf$ has local maxima and there are $N\geq 1$ substantial among them, say $x_{1}<x_{2}<\dots<x_{N}$, then for $N\geq 2$ and every $2\leq n\leq N$ we choose a local minimum $y_{n}\in(x_{n-1},x_{n})$. Then
${\rm Var}(Mf)={\rm Var}_{(-\infty,x_{1}]}(Mf)+\Big{(}\sum_{n=2}^{N}Mf(x_{n-1})-2Mf(y_{n})+Mf(x_{n})\Big{)}+{\rm Var}_{[x_{N},\infty)}(Mf),$
with the middle term disappearing when $N=1$.
Let $T_{I}(x)\coloneqq\frac{|I|}{|J_{I}(x)|}$ for all $x\in\mathbb{R}$, where $J_{I}(x)$ is the smallest interval $J\supseteq I$ with $c(J)=x$. The next result says that the variation of $Mf$ is controlled locally by the variation of $T_{I_{k}}$, $1\leq k\leq K$.
Lemma 3.2.
Suppose $Mf$ has $N\geq 2$ substantial local maxima $x_{1}<\dots<x_{N}$. Then
$Mf(x_{n-1})-Mf(y_{n})\leq\sum_{k=1}^{K}\alpha_{k}{\rm Var}_{[x_{n-1},y_{n}]}(T_{I_{k}})\quad\text{and}\quad Mf(x_{n})-Mf(y_{n})\leq\sum_{k=1}^{K}\alpha_{k}{\rm Var}_{[y_{n},x_{n}]}(T_{I_{k}})$
for all $2\leq n\leq N$ and $x_{1}<y_{2}<x_{2}<\dots<x_{N-1}<y_{N}<x_{N}$. Moreover, if $l(I_{1})=A$ and $\alpha_{1}\leq a$, then the first summand in each sum above may be omitted (note that $N\geq 2$ implies $K\geq 2$).
Proof.
We shall prove the second inequality, the proof of the first one is analogous.
Given $x_{n}$, we take $I_{x_{n}}$ as in the conclusions of Lemma 3.1 with accompanying $1\leq k^{\prime}\leq k^{\prime\prime}\leq K$. Then, for $I_{y_{n}}$ centered at $y_{n}$ and such that $r(I_{y_{n}})=r(I_{x_{n}})$, we have
$Mf(x_{n})-Mf(y_{n})\leq\frac{1}{|I_{x_{n}}|}\int_{I_{x_{n}}}f-\frac{1}{|I_{y_{n}}|}\int_{I_{y_{n}}}f\leq\sum_{k=k^{\prime}}^{k^{\prime\prime}}\frac{\alpha_{k}|I_{k}|}{|I_{x_{n}}|}-\frac{\alpha_{k}|I_{k}|}{|I_{y_{n}}|}.$
Observe that $k^{\prime}=1$ cannot happen if $l(I_{1})=A$ and $\alpha_{1}\leq a$ because then $\alpha_{1}\geq Mf(x_{n})$ implies $Mf(y)\geq Mf(x_{n})$ for all $y\leq x_{n}$; this explains the last sentence of the statement. Next, we notice that the numerical inequality $\frac{1}{\rho_{2}}-\frac{1}{\rho_{2}+\rho_{0}}\leq\frac{1}{\rho_{1}}-\frac{1}{\rho_{1}+\rho_{0}}$ holds for $\rho_{0},\rho_{1},\rho_{2}\in(0,\infty)$ when $\rho_{1}\leq\rho_{2}$. Thus, if either $x_{n}\leq c(I_{k})$ or $c(I_{k})\leq y_{n}$, then
$\frac{\alpha_{k}|I_{k}|}{|I_{x_{n}}|}-\frac{\alpha_{k}|I_{k}|}{|I_{y_{n}}|}\leq\alpha_{k}\Big{|}\frac{|I_{k}|}{|J_{I_{k}}(x_{n})|}-\frac{|I_{k}|}{|J_{I_{k}}(y_{n})|}\Big{|}=\alpha_{k}|T_{I_{k}}(x_{n})-T_{I_{k}}(y_{n})|=\alpha_{k}{\rm Var}_{[y_{n},x_{n}]}(T_{I_{k}}).$
In the remaining case, $y_{n}<c(I_{k})<x_{n}$, we add and subtract $\frac{\alpha_{k}|I_{k}|}{|I^{\prime}|}$ with $I^{\prime}$ such that $c(I^{\prime})=c(I_{k})$ and $r(I^{\prime})=r(I_{y_{n}})=r(I_{x_{n}})$, which gives
$\alpha_{k}\bigg{(}\frac{|I_{k}|}{|I_{x_{n}}|}-\frac{|I_{k}|}{|I^{\prime}|}+\frac{|I_{k}|}{|I^{\prime}|}-\frac{|I_{k}|}{|I_{y_{n}}|}\bigg{)}\leq\alpha_{k}\bigg{(}\Big{|}\frac{|I_{k}|}{|J_{I_{k}}(x_{n})|}-\frac{|I_{k}|}{|J_{I_{k}}(c(I_{k}))|}\Big{|}+\Big{|}\frac{|I_{k}|}{|J_{I_{k}}(c(I_{k}))|}-\frac{|I_{k}|}{|J_{I_{k}}(y_{n})|}\Big{|}\bigg{)}$
and the last quantity does not exceed $\alpha_{k}\big{(}2T_{I_{k}}(c(I_{k}))-T_{I_{k}}(x_{n})-T_{I_{k}}(y_{n})\big{)}=\alpha_{k}{\rm Var}_{[y_{n},x_{n}]}(T_{I_{k}})$. ∎
Proof of Theorem 1.1.
For simplicity, we assume $a\geq b$. If $Mf$ has no local maxima, then
${\rm Var}(Mf)=\frac{a-b}{2}=\big{(}f(+\infty)-f(-\infty)\big{)}-\frac{a-b}{2}\leq{\rm Var}(f)-\frac{a-b}{2}.$
The same conclusion easily follows when $Mf$ has exactly one substantial local maximum. Assume therefore that $N\geq 2$ and $x_{1}<\dots<x_{N}$ are all substantial local maxima of $Mf$. Since $Mf\geq\frac{a+b}{2}$, $T_{I_{k}}(\pm\infty)=0$, and $Mf$ decreases on $[x_{N},\infty)$, arguing as in the proof of Lemma 3.2 gives
${\rm Var}_{[x_{N},\infty)}(Mf)=Mf(x_{N})-\frac{a+b}{2}\leq\sum_{k=1}^{K}\alpha_{k}{\rm Var}_{[x_{N},\infty)}(T_{I_{k}})-\frac{a+b}{2},$
with the first summand omitted if $l(I_{1})=A$ and $\alpha_{1}\leq a$. If $Mf$ is nondecreasing on $(-\infty,x_{1}]$, then
${\rm Var}_{(-\infty,x_{1}]}(Mf)=Mf(x_{1})-a\leq\sum_{k=1}^{K}\alpha_{k}{\rm Var}_{(-\infty,x_{1}]}(T_{I_{k}})-a.$
In this case combining the two estimates above with these from Lemma 3.2 gives (1.2), since
${\rm Var}(Mf)\leq\sum_{k=1}^{K}\alpha_{k}{\rm Var}(T_{I_{k}})-a-\frac{a+b}{2}=\sum_{k=1}^{K}2\alpha_{k}-a-b-\frac{a-b}{2}\leq{\rm Var}(f)-\frac{a-b}{2}.$
Otherwise, if $x_{1}$ is the first local maximum of $Mf$ but $Mf$ is not nondecreasing on $(-\infty,x_{1}]$, then there is a local minimum $y_{1}\in[A,x_{1})$ such that
${\rm Var}_{(-\infty,x_{1}]}(Mf)=a-2Mf(y_{1})+Mf(x_{1})\leq\min\Big{\{}Mf(x_{1})-b,a-\frac{a+b}{2}+\sum_{k=1}^{K}\alpha_{k}{\rm Var}_{[y_{1},x_{1}]}(T_{I_{k}})\Big{\}}.$
The above inequality, which splits into two parts, is a consequence of $(a+b)/2\leq Mf$ (first part), and the argument used in the proof of Lemma 3.2 (second part) with the additional nonnegative term $Mf(y_{1})-(a+b)/2$ added. Again the first summand is not used if $l(I_{1})=A$ and $\alpha_{1}\leq a$.
We now analyze separately three cases depending on the relation between $A$, $l(I_{1})$, $a$, and $\alpha_{1}$. If $l(I_{1})>A$, then we choose the second term from the above minimum and obtain (1.2), since after splitting ${\rm Var}(Mf)$ with respect to $(-\infty,x_{1}]$, $[x_{n},x_{n+1}]$, $n=1,\dots,N-1$, and $[x_{N},\infty)$, we obtain
${\rm Var}(Mf)\leq a-\frac{a+b}{2}+\sum_{k=1}^{K}\alpha_{k}{\rm Var}(T_{I_{k}})-\frac{a+b}{2}=\sum_{k=1}^{K}2\alpha_{k}-b\leq{\rm Var}(f)-a.$
If $l(I_{1})=A$ and $\alpha_{1}\leq a$, then ${\rm Var}(T_{I_{1}})$ is not used. Thus, choosing again the second term, we have
${\rm Var}(Mf)\leq\sum_{k=2}^{K}2\alpha_{k}-b=a-\alpha_{1}+\alpha_{1}+\sum_{k=2}^{K}2\alpha_{k}-b-a\leq{\rm Var}(f)-a.$
Finally, if $l(I_{1})=A$ and $\alpha_{1}>a$, then we choose the first term from the relevant minimum. Now, $x_{1}>r(I_{1})$ because otherwise we would have $Mf(y_{1})>a$. Moreover, $Mf(x_{1})\geq\alpha_{1}$ because $x_{1}$ is the first local maximum, while $Mf(x_{1})=\alpha_{1}$ cannot happen because then $Mf(y)=\alpha_{1}$ for all $y\in[c(I_{1}),x_{1}]$ and again $Mf(y_{1})>a$. Consequently, $l(I_{x_{1}})=l(I_{k^{\prime}})$ for some $k^{\prime}>1$ with $\alpha_{k^{\prime}}>\alpha_{1}>a$, and we have $x_{1}\geq c(I_{k^{\prime}})$. Thus, we obtain
${\rm Var}(Mf)\leq\alpha_{k^{\prime}}-b+\sum_{k=1}^{K}\alpha_{k}{\rm Var}_{[x_{1},\infty)}(T_{I_{k}})-\frac{a+b}{2}\leq(a-\alpha_{1})+{\rm Var}(f)-\frac{a+b}{2}<{\rm Var}(f)-\frac{a-b}{2},$
since by $x_{1}\geq c(I_{k^{\prime}})\geq c(I_{1})$ the above sum does not exceed $2(\alpha_{1}+\dots+\alpha_{K})-\alpha_{1}-\alpha_{k^{\prime}}$. ∎
4. Further comments
Here we collect several thoughts that complement our main results. We verify optimality of the constants in Conjecture $\Delta$, discuss variants of (1.2) and (1.3) for other operators, comment on the special case of indicator functions, and point out limitations of our proof method.
First, we note that the constants $1$ and $\frac{1}{2}$ in (1.1) are optimal. Indeed, for $a,b\in[0,\infty)$ with $a\geq b$ and $N\in\mathbb{N}$, take $g_{N}\in{\rm BV}_{a}^{b}(\mathbb{R})$ defined by
$g_{N}\coloneqq a{\mathds{1}_{{(-\infty,-1]}}}+N{\mathds{1}_{{(-1,1)}}}+b{\mathds{1}_{{[1,\infty)}}}.$
Then for $N\geq a$, a simple calculation shows ${\rm Var}(g_{N})=2N-a-b$ and ${\rm Var}(Mg_{N})=2N-a-\frac{a+b}{2}$. Letting $N\to\infty$, we obtain $c(\mathbb{R})\geq 1$ and $C_{a}^{b}(\mathbb{R})\leq\frac{a-b}{2}$.
Next, we show that certain variants of Conjecture $\Delta$ and its discrete counterpart hold for other maximal operators. Set
$\overrightarrow{M}g(x)\coloneqq\sup_{r>0}\frac{1}{r}\int_{x}^{x+r}|g|\quad\text{and}\quad\overrightarrow{M}G(n)\coloneqq\sup_{m\geq 0}\frac{1}{m+1}\sum_{y=n}^{n+m}|G(y)|,$
and, similarly,
$\widetilde{M}g(x)\coloneqq\sup_{r_{1},r_{2}>0}\frac{1}{r_{1}+r_{2}}\int_{x-r_{1}}^{x+r_{2}}|g|\quad\text{and}\quad\widetilde{M}G(n)\coloneqq\sup_{m_{1},m_{2}\geq 0}\frac{1}{m_{1}+m_{2}+1}\sum_{y=n-m_{1}}^{n+m_{2}}|G(y)|.$
We call $\overrightarrow{M}$ and $\widetilde{M}$ the one-sided and uncentered maximal operators, respectively. When $\mathbb{K}=\mathbb{Z}$, one can instead consider a variant of $\widetilde{M}$ allowing only $m_{1},m_{2}\geq 0$ with $m_{1}+m_{2}+1\in 2\mathbb{N}-1$ so that the corresponding intervals are balls in $\mathbb{Z}$. This does not change the conclusion below.
Since the sampling procedure works in the context of $\overrightarrow{M}$ or $\widetilde{M}$ with only minor adjustments, for all $a,b\in\mathbb{C}$ the associated constants $\overrightarrow{c}(\mathbb{R}),\overrightarrow{C_{a}^{b}}(\mathbb{R}),\widetilde{c}(\mathbb{R}),\widetilde{C_{a}^{b}}(\mathbb{R})$ are not worse than their discrete counterparts. Furthermore, when $\mathbb{K}=\mathbb{Z}$ and $a,b\in[0,\infty)$, it is natural to conjecture that $\overrightarrow{c}(\mathbb{Z})=\widetilde{c}(\mathbb{Z})=1$, $\overrightarrow{C_{a}^{b}}(\mathbb{Z})=\max\{0,b-a\}$, and $\widetilde{C_{a}^{b}}(\mathbb{Z})=|b-a|$. Indeed, a suitable variant of (1.3) follows because for each nonnegative $G\in{\rm BV}_{a}^{b}(\mathbb{Z})$ one has $\overrightarrow{M}G\geq\max\{G,b\}$ and $\widetilde{M}G\geq\max\{G,a,b\}$, while the equality $\overrightarrow{M}G(n)=G(n)$ or $\widetilde{M}G(n)=G(n)$ holds whenever $n$ is a local maximum of $\overrightarrow{M}G$ or $\widetilde{M}G$. Considering $G_{N}\in{\rm BV}_{a}^{b}(\mathbb{Z})$ given by $G_{N}\coloneqq a{\mathds{1}_{{-\mathbb{N}}}}+N{\mathds{1}_{{\{0\}}}}+b{\mathds{1}_{{\mathbb{N}}}}$ and $g_{N}\in{\rm BV}_{a}^{b}(\mathbb{R})$ given by $g_{N}\coloneqq\sum_{n\in\mathbb{Z}}G_{N}(n){\mathds{1}_{{[n-1/2,n+1/2)}}}$, we verify that the postulated constants are optimal.
Corollary 4.1.
For all $a,b\in[0,\infty)$ we have
  1. (1)
    $\overrightarrow{c}(\mathbb{R})=\overrightarrow{c}(\mathbb{Z})=1$ and $\overrightarrow{C_{a}^{b}}(\mathbb{R})=\overrightarrow{C_{a}^{b}}(\mathbb{Z})=\max\{0,b-a\}$,
  2. (2)
    $\widetilde{c}(\mathbb{R})=\widetilde{c}(\mathbb{Z})=1$ and $\widetilde{C_{a}^{b}}(\mathbb{R})=\widetilde{C_{a}^{b}}(\mathbb{Z})=|b-a|$.
In the case of indicator functions our proof of (1.2) and (1.3) is simpler. Indeed, let $f\in{\rm BV}(\mathbb{R})$ be as in Theorem 1.1 with $a,b,\alpha_{k}\in\{0,1\}$. Now, if $x$ is a local maximum of $Mf$, then either $f(x)=0$ or $x$ can be adjusted to $x_{*}=c(I_{k})$ for some $k$. On the other hand, if $y$ is a local minimum, then $f(y)=0$. Thus, we can replace $T_{I_{k}}$ by $M{\mathds{1}_{{I_{k}}}}$, since both functions coincide on $\{c(I_{k})\}\cup(\mathbb{R}\setminus I_{k})$. In particular, for each closed interval $J$ whose endpoints are neighboring extrema $x,y$ we obtain
${\rm Var}_{J}(Mf)=Mf(x)-Mf(y)\leq\frac{1}{|I_{x}|}\int_{I_{x}}|f|-\frac{1}{|I_{y}|}\int_{I_{y}}|f|\leq\sum_{k=1}^{K}\alpha_{k}{\rm Var}_{J}(M{\mathds{1}_{{I_{k}}}}).$
[…]
Finally, notice that the method presented in the proof of Theorem 1.1 has some natural limitations. Indeed, during the proof we always choose $I_{y}$ to be the smallest interval centered at $y$ which contains $I_{x}$. This choice is sometimes very far from optimal. To illustrate this obstacle, let us take $g\in{\rm BV}_{0}^{0}(\mathbb{R})$ given by $g\coloneqq{\mathds{1}_{{[-2,-1]\cup[1,2]}}}+C{\mathds{1}_{{[-2,2]}}}$ with $C=10^{6}$, say. Then $Mg$ has local maxima at $-\frac{3}{2},0,\frac{3}{2}$ with the related values being $C+1,C+\frac{1}{2},C+1$, and local minima at $-\frac{1}{2},\frac{1}{2}$ with the related values being $C+\frac{1}{3},C+\frac{1}{3}$. Thus, ${\rm Var}(Mg)=2C+4-\frac{1}{3}\leq 2C+4={\rm Var}(g)$, as desired. However, for $x=0$ we have $I_{x}=(-2,2)$ so that for $y=\frac{1}{2}$ we should take $I_{y}=(-2,3)$ in Theorem 1.1. Unfortunately, the difference between the corresponding averages is much larger than $Mg(0)-Mg(\frac{1}{2})$, therefore ${\rm Var}(Mg)\leq{\rm Var}(g)$ cannot be verified this way.
Acknowledgments
Paul Hagelstein is partially supported by a grant from the Simons Foundation (#521719 to Paul Hagelstein) and Dariusz Kosz is partially supported by a grant from the National Science Centre of Poland (SONATA BIS 2022/46/E/ST1/00036).